decide whether the pair of lines is , parallel, perpendicular, or none....
4x+3y=10
3x+4y=2

Answer by stanbon(75887)

decide whether the pair of lines is , parallel, perpendicular, or none....
4x+3y=10
3x+4y=2
---
Check the slopes:
1st: slope = -4/3
2nd: slope = -3/4
-------------------------------------------------
Ans: none; they are not perpendicular and not parallel
==================
Cheers,
Stan H.